1.  Given that z represents a standard normal random variable, provide the probabilities for the following:

  1. P (z < 1.56)      _____________

  1. P (z > -.94)      _____________

  1. P (1.57 < z < 2.30)   ____________
Answer #1

solution

(A)P(z <1.56 )

Using z table  

=0.9406

probability =0.9406

(B)

P(z >-0.94 ) =1 - P(z <-0.94 )

Using z table,
= 1 -0.1736

=0.8264

probability = 0.8264   

(C)P (1.57 < z < 2.30)

= P(Z < 2.30) - P(Z <1.57 )

Using z table,  

=   0.9893-0.9418

probability=0.0475